OpenStudy (anonymous):

Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). (5 points) Part B: How much time does Ted spend on practicing the breast stroke every day? (3 points) Part C: Is it possible for Ted to have spent 45 minutes practicing the butterfly stroke every day?

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

He practices for 50 min each day: x + y = 50 And does 20 min more of breast stroke than the other: y = x + 20 There are your two equations for Part A. --- for part B, solve for x and y. The specific answer they are looking for is "y", so I will solve one equation for "x" and then substitute that expression in terms of y into the other, then solve for y: x + y = 50 x = 50 - y y = x + 20 y = 50 - y + 20 2y = 70 y = 35 He spends 35 min each day doing breast stroke, and therefore 15 min doing butterfly stroke. --- For part C, I'm not sure what the solution is after. Is it possible for him to practice butterfly for 45 min every day? Sure. but if he wants to limit to 50 min, he has to change the time practicing the other. Or if he still wants to practice the breast stroke for 20 min more than butterfly, he'll have to allot more than 50 min per day for swimming practice. Something has to change.

OpenStudy (anonymous):

i did it on my own i just wanna make sure its right

imqwerty (imqwerty):

http://prntscr.com/8po3w6 XD

imqwerty (imqwerty):

np and i think that 3rd one should be yes and for that he must change his schedule :D

OpenStudy (anonymous):

ok ty ty

imqwerty (imqwerty):

np (: